OpenStudy (stylesprettygirl):

Can anyone help me with simple interest Im still not 100 % sure how to solve these problems any help would be greatly apperciated

OpenStudy (stylesprettygirl):

Thomas opened a savings account with an annual interest rate of 7% and an initial deposit of $5000. If his interest is compounded quarterly, how much is in Thomas’s account after 4 years? Round your answer to the nearest cent. interest compounded quarterly: A = P (1 +fraction numerator r over denominator 4 end fraction)4t

OpenStudy (stylesprettygirl):

A = 5000(1+7/4)^4x4 is this right and if it is how do I get the answer from here?

OpenStudy (pyxieyt):

Do you know the formula for simple interest?

OpenStudy (pyxieyt):

You filled it in properly!

OpenStudy (stylesprettygirl):

Yes but its interest compound quartetly

OpenStudy (pyxieyt):

You know pemdas right? or order of operations

OpenStudy (pyxieyt):

Here it is if you need it

OpenStudy (stylesprettygirl):

I know pemdas not order of operations tho

OpenStudy (pyxieyt):

Pemdas is the order of operations :D

OpenStudy (stylesprettygirl):

ohhhh so () ^ x division + -

OpenStudy (pyxieyt):

Yep!

OpenStudy (pyxieyt):

So with A = 5000(1+7/4)^4x4 what do you think you start with?

OpenStudy (stylesprettygirl):

so the answer would be 6599.65

OpenStudy (stylesprettygirl):

am I right?

OpenStudy (pyxieyt):

Let me check again

OpenStudy (pyxieyt):

you double check too

OpenStudy (stylesprettygirl):

is it right?

OpenStudy (pyxieyt):

This is not multiple choice right?

OpenStudy (stylesprettygirl):

nope

OpenStudy (pyxieyt):

Okay just had to make sure xD

OpenStudy (stylesprettygirl):

which answer did you get

OpenStudy (pyxieyt):

Something different but i shall check one more time xD

OpenStudy (stylesprettygirl):

what?

OpenStudy (pyxieyt):

Nothing but i think you are right but I recommend getting help! ( i need to leave for a second they have a bug on my computer xD)